Find the radius and the centre of the circular section of the sphere |r|= 4, cut off by the plane r.(2i-j+4k)=3.


1
Expert's answer
2021-05-24T11:52:38-0400


Let "S" be the sphere in "R^3" with center "O(x_0,y_0,z_0)" and radius "r," and let "P" be the plane with equation "Ax+By+Cz+D=0," so that "\vec n=(A,B,C)"is a normal vector of "P."

If "P_0" is an arbitrary point on "P," the signed distance from the center of the sphere "O" to the plane "P" is


"\rho=\dfrac{(O-P_0)\vec n}{\|\vec n\|}=\dfrac{Ax_0+By_0+Cz_0-D}{\sqrt{A^2+B^2+C^2}}."


The intersection "S\cap P" is a circle if and only if "r<\rho<r," and in that case, the circle has radius "r_c=\sqrt{r^2-\rho^2}" and center


"C=O+\rho \cdot\dfrac{\vec n}{\|\vec n\|}"

"=(x_0, y_0, z_0)+\rho\cdot\dfrac{(A,B,C)}{\sqrt{A^2+B^2+C^2}}"


In our case:

"O(0,0,0), |r|=4, 2x-y+4z=3"


"S=\{(x,y,z):x^2+y^2+z^2=16\},"


"P=\{(x,y,z):2x-y+4z=3\}\}."



"\rho=\dfrac{Ax_0+By_0+Cz_0-D}{\sqrt{A^2+B^2+C^2}}"

"=\dfrac{2\cdot0-1\cdot0+4\cdot0-3}{\sqrt{(2)^2+(-1)^2+(4)^2}}=-\dfrac{3}{\sqrt{21}}"

"r_c=\sqrt{r^2-\rho^2}=\sqrt{(4)^2-(-\dfrac{3}{\sqrt{21}})^2}=\sqrt{\dfrac{109}{7}}"

"\approx3.946"


"C=O+\rho \cdot\dfrac{\vec n}{\|\vec n\|}"

"=(0, 0, 0)+\big(-\dfrac{3}{\sqrt{21}}\big)\cdot\dfrac{(2,-1,4)}{\sqrt{(2)^2+(-1)^2+(4)^2}}"

"=(-\dfrac{2}{7}, \dfrac{1}{7}, -\dfrac{4}{7})"


"r_c=\sqrt{\dfrac{109}{7}}\approx3.946, C=(-\dfrac{2}{7}, \dfrac{1}{7}, -\dfrac{4}{7})"